# Localize-and-Stitch: Efficient Model Merging via Sparse Task Arithmetic

### v1

@@ 5 Related works
Model merging. Model merging aims at efficiently integrating multiple finetuned models into a single model that retains the capabilities of each. It is a promising solution to enhance the generalization and multi-task capabilities of finetuned models. Wortsman et al. [65, 64] demonstrates that merging models trained on the same task with different training configurations can improve the out-of-distribution (OOD) robustness. Merging finetuned models from related tasks can provide an improved initialization for new tasks [7, 20]. Additionally, finetuned models with different specialized capabilities can be merged together for enhanced multi-task capabilities [30, 67, 40, 32, 68, 70, 60]. In this work, we primarily focus on merging model for enhancing multi-task performance. Similar to the gradient conflict problem [71, 37] in multi-task learning, finetuned models also manifest conflict when merged together, and our method provides an effective solution to this problem.
@@
Knowledge attribution. Recent works find that knowledge contained in language models is localizable, meaning that model behavior can be attributed to small subsets of model parameters. One line of work identifies such regions to edit the knowledge contained in the networks. Sundararajan et al. [55] uses integrated gradients for knowledge attribution, which measures how sensitive each neuron’s gradient is to the change of input. Dai et al. [12] applies integrated gradients to edit factual knowledge contained in BERT models. However, the relationship between the editing success and the localized regions remains unclear [22]. Another line of work aims at pruning large models. Lee et al. [35] proposes SNIP score, which computes the change of loss when each neuron is set to 0. Sun et al. [54] proposes Wanda score to optimize for a binary mask such that the masked activations is close to the original ones. Knowledge attribution can also be applied for enhancing interpretability. Vig et al. [57] applies causal mediation analysis [46] to identify individual neurons contributing to gender bias.
Recently, Panigrahi et al. [45] optimizes for a binary mask to localize the skills contained in finetuned language models and study task relationship based on it. There are two key differences between our localization formulation and theirs. Firstly, our formulation of $S$ is more straightforward, as we directly have $\gamma=\sigma(S)$ in Equation 1. In contrast, Panigrahi et al. [45] uses $S$ as a selector of whether to take the value from the initial mask, leading to more complex computation. Secondly, our approach uses the $L_{1}$ constraint to control the sparsity in a more fine-grained manner, while Panigrahi et al. [45] does not have this constraint, and controls sparsity via early stopping. We empirically show that our localization formulation identifies regions with improved quality in Appendix C.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
Han Zhao and Tong Zhang are partially supported by an NSF IIS grant No. 2416897. Han Zhao would also like to thank the support from a Google Research Scholar Award. Yifei He and Yuzheng Hu are partially supported by a research grant from the Amazon-Illinois Center on AI for Interactive Conversational Experiences (AICE). The views and conclusions expressed in this paper are solely those of the authors and do not necessarily reflect the official policies or positions of the supporting companies and government agencies.
Appendix A More experiments
@@
Where are the localized regions? We analyze the distribution of the localized regions for both language and vision tasks in Figure 9, both in terms of the layer index and the transformer components. For the layers, different tasks seem to occupy different layers, although the earlier layers in the network seldomly appear in the localized regions. Interestingly, most of the localized regions concentrate in the LayerNorm parameters. This pattern can possibly be attributed to a covariate shift observed in the finetuning data compared to the pretraining data, necessitating adjustments to the LayerNorm parameters to accommodate this shift.
Table 5: Single-task grafted performance of RoBERTa-base models on twelve NLP tasks.
Task SST-2 CR MR MPQA TREC SUBJ QNLI SNLI MNLI RTE MRPC QQP Average
@@
Table 6: Single-task grafted performance of GPT2-XL models on three NLP tasks.
Task MMLU TruthfulQA ARC Average
@@
Table 7: Single-task grafted performance of CLIP ViT-B/32 models on eight vision tasks.
Task SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Average
@@
Full grafted performance. We evaluate the quality of the localized regions by the grafted performance. For the $i$-th task, we only add up the localized regions in the task vectors back to the pretrained model, i.e., $\theta_{\text{pre}}+\gamma_{i}\odot\tau_{i}$. The results with a localization region of $1\%$ is shown in Tables 5, 6 and 7. For almost all tasks, using only the tiny localized region recovers nearly $99\%$ of the finetuned performance. For GPT2-XL, the performance is slightly worse because we cannot use the evaluation data for the localization step. However, the results are still strong even with surrogate datasets with similar purposes, demonstrating the flexibility and robustness of our algorithm. Overall, this shows that our localization approach is effective in locating regions containing essential skills acquired during finetuning, and the localized regions can be viewed as compact representations of the finetuned models.
@@
Runtime. We compare the runtime of Localize-and-Stitch and other baselines. We divide the algorithms into two categories: dataless and requiring data. Note that task arithmetic and TIES-Merging can fall in both categories, with the difference of whether performing hyperparameter tuning (scaling factor $\alpha$ for both, and sparsity for TIES). For the hyperparameter tuning, we follow the common practice in [30, 67] to grid search over $\{0.1,0.2\cdots,0.9\}$ for the scaling factor and $\{0.1,0.2,0.3\}$ for the sparsity.
We report the runtime in Tables 9 and 9 for merging twelve NLP tasks with RoBERTa-base. For the dataless algorithms, simple averaging and task arithmetic are very efficient, as they only involve arithmetic operations on the weights. Both TIES and our dataless version requires sorting the task vectors to get the top-$k\%$ largest elements, so the runtime is slower. Compared with TIES, we do not have the step for resolving sign conflicts, so it takes less time. For algorithms requiring data, Fisher merging is the most efficient, as it uses a diagonal estimate of the Fisher information matrices with little data (256 per task). Both task arithmetic and TIES-Merging show substantial time increase, as they need to do grid search on 9 and 27 hyperparameters respectively as well as evaluating on the validation data in each run. AdaMerging takes significantly more runtime to execute compared with others, and the reason could be that entropy minimization converges slowly, as we observe that AdaMerging requires around 500 epochs to converge. Compared with other algorithms, Localize-and-Stitch executes in a relatively short amount of time, showing its effectiveness.
Table 10: Comparison of performance and efficiency for continual learning.
SST-2+CR +TREC +SUBJ +QNLI +QQP
Task arithmetic Average performance 0.907 0.864 0.853 0.817 0.762
Runtime (s) 1340 1760 2183 2654 2970
TIES Average performance 0.897 0.832 0.827 0.796 0.757
Runtime (s) 3421 5031 6894 8147 9358
Localize-and-Stitch Average performance 0.906 0.905 0.897 0.856 0.827
Runtime (s) 806 421 403 462 455
Continual learning. As mentioned in Section 3.3, our approach is particularly efficient in the continual learning setting. To illustrate the efficiency, we start from merging the SST-2 and CR RoBERTa models, and incrementally add 4 more tasks to simulate a continual learning setting. The tasks are selected by representativeness (including sentiment analysis, sentence classification, NLI etc). We compare our method against two baselines: task arithmetic and TIES, with merging coefficients tuned across $\{0.0,0.1,...0.9,1.0\}$. The results from Table 10 demonstrate that: i) Performance: Localize-and-Stitch consistently outperforms the baselines, with the performance margin increasing as more tasks are involved, showing its ability to reduce task interference. ii) Runtime: The runtime for both TIES and task arithmetic increases with each added task due to the need for hyperparameter search from scratch and performance validations, while the runtime of Localize-and-Stitch generally remains constant, reflecting its efficiency in continual learning scenarios.
Appendix B Comparison between Dataless Localize-and-Stitch and TIES-Merging
Due to the similarity of the dataless version of our approach with TIES-Merging [67], we compare them in detail.
Table 11: Comparison of Dataless Localize-and-Stitch and TIES on language tasks.
Task SST-2 CR MR MPQA TREC SUBJ QNLI SNLI MNLI RTE MRPC QQP Average
@@ Appendix B Comparison between Dataless Localize-and-Stitch and TIES-Merging
Table 12: Comparison of Dataless Localize-and-Stitch and TIES on vision tasks.
Task SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Average
@@
The first step of both algorithms is similar: select the top-$k\%$ largest positions in the task vector. The primary difference lies in the selection threshold: Dataless Localize-and-Stitch selects the top-$5\%$, while TIES selects the top-$20\%$. However, as shown in Section 3.1, we find that when a localized region already contains sufficient task-specific knowledge, including more parameters only introduces more task interference. This observation could partially explain our superior performance. However, this is not the only limitation in TIES, as reducing the threshold in TIES to be $5\%$ does not yield an improved performance as demonstrated in Tables 11 and 12.
@@
Appendix C Details on localization
Here, we detail the skill attribution method in [45] and explain the difference with our formulation. Panigrahi et al. [45] aims to localize task-specific skills contained in finetuned language models. They introduce model grafting, where for given pretrained and finetuned model parameters $\theta_{\text{pre}}$ and $\theta_{\text{ft}}$, they graft parameters of $\theta_{\text{ft}}$ in the region $\gamma$ onto the pretrained model as

### v2

@@ 5 Related works
Model merging. Model merging aims at efficiently integrating multiple finetuned models into a single model that retains the capabilities of each. This approach enhances the efficiency, generalization and multi-task capabilities of finetuned models. In scenarios where models are trained on the same task with different training configurations, Singh & Jaggi (2020); Ainsworth et al. (2022); Jolicoeur-Martineau et al. (2024) show that merged models perform comparably to ensemble models but with significantly lower deployment costs. Additionally, Wortsman et al. (2022a; b) demonstrate that the merged model improves the out-of-distribution (OOD) robustness. When merging finetuned models from different tasks, the merged model can provide better initialization for new tasks (Choshen et al., 2022; Gueta et al., 2023). Finetuned models with different specialized skills can also be combined to enhance multi-task capabilities (Ilharco et al., 2023; Tam et al., 2023; Matena & Raffel, 2022; Jin et al., 2022; Yang et al., 2023; Yu et al., 2023; Wang et al., 2024b; a). More recently, a new line of work has emerged that uses a mixture of experts (MoE) strategy (Jiang et al., 2023; Tang et al., 2024). Instead of a single unified model, the MoE approach incorporates routing mechanisms to direct inputs to task-specific networks. In this work, we primarily focus on merging specialized models into a single unified model for enhancing multi-task performance. Similar to the gradient conflict problem (Yu et al., 2020; Liu et al., 2021) in multi-task learning, finetuned models also manifest conflict when merged together, and our method provides an effective solution to this problem.
@@
Knowledge attribution. Recent works find that knowledge contained in language models is localizable, meaning that model behavior can be attributed to small subsets of model parameters. One line of work identifies such regions to edit the knowledge contained in the networks. Sundararajan et al. (2017) uses integrated gradients for knowledge attribution, which measures how sensitive each neuron’s gradient is to the change of input. Dai et al. (2021) applies integrated gradients to edit factual knowledge contained in BERT models. However, the relationship between the editing success and the localized regions remains unclear (Hase et al., 2024). Knowledge attribution can also be applied for enhancing interpretability. Vig et al. (2020) applies causal mediation analysis (Pearl, 2022) to identify individual neurons contributing to gender bias.
Recently, Panigrahi et al. (2023) optimizes for a binary mask to localize skills contained in finetuned language models. There are two key differences between our localization formulation and theirs. Firstly, our formulation of $S$ is more straightforward, as we directly have $\gamma=\sigma(S)$ in Equation 1. In contrast, Panigrahi et al. (2023) uses $S$ as a selector of whether to take the value from the initial mask, leading to more complex computation. Secondly, our approach uses the $L_{1}$ constraint to control the sparsity in a more fine-grained manner, while Panigrahi et al. (2023) does not have this constraint, and controls sparsity via early stopping. We empirically show that our localization formulation identifies regions with improved quality in Appendix E.
Pruning. Similar to localization, pruning is a strategy to identify key regions in the parameter space that are important to model performance. Han et al. (2015a; b) propose magnitude pruning, which preserves weights with high magnitudes, and the method inspires various variants (Zhu & Gupta, 2017; Paganini & Forde, 2020; Zafrir et al., 2021). Parameter significance measured by performance sensitivity is another effective criteria for identifying important parameters (Sanh et al., 2020; Liang et al., 2021; Zhang et al., 2022). For instance, Lee et al. (2018) proposes SNIP score, which computes the change of loss when each neuron is set to 0. Pruning methods are widely applied to modern large language models (Zhang et al., 2023; Sun et al., 2023; Xia et al., 2023; Zhao et al., 2024; Cheng et al., 2024).
The primary distinction between pruning and localization lies in their treatment of parameters outside the identified regions. In pruning, these parameters are set to zero, effectively removing them, allowing the pruned network to function as a standalone model. Conversely, in localization, parameters outside the localized regions are retained at their pretrained values, requiring the localized regions to be combined with the pretrained model for deployment. Despite these differences, the conceptual overlap between them suggests that pruning techniques could be adapted for localization. Exploring such adaptations presents an interesting direction for future work.
@@ 6 Conclusion
Appendix A Full experimental results
Table 4: Multi-task accuracy (F1) of merged RoBERTa-base models on twelve NLP tasks.
Task SST-2 CR MR MPQA TREC SUBJ QNLI SNLI MNLI RTE MRPC QQP Average Acc/F1
Simple averaging (Wortsman et al., 2022a) 0.857 0.851 0.829 0.688 0.304 0.478 0.508 0.452 0.452 0.563 0.311 0.469 0.563
Task arithmetic (Ilharco et al., 2023) 0.846 0.856 0.769 0.810 0.156 0.584 0.607 0.538 0.403 0.539 0.822 0.581 0.626
TIES (Yadav et al., 2023) 0.805 0.805 0.728 0.791 0.226 0.549 0.552 0.501 0.379 0.477 0.816 0.572 0.600
Dataless Localize-and-Stitch 0.909 0.907 0.864 0.821 0.462 0.762 0.558 0.690 0.618 0.688 0.837 0.693 0.734
Task arithmetic (Ilharco et al., 2023) 0.885 0.882 0.803 0.829 0.320 0.610 0.620 0.561 0.495 0.656 0.828 0.623 0.675
TIES (Yadav et al., 2023) 0.886 0.88 0.852 0.835 0.226 0.482 0.548 0.359 0.397 0.594 0.794 0.603 0.621
Fisher merging (Matena & Raffel, 2022) 0.900 0.898 0.837 0.758 0.260 0.546 0.542 0.725 0.652 0.656 0.833 0.677 0.690
RegMean (Jin et al., 2022) 0.897 0.897 0.847 0.826 0.730 0.791 0.559 0.683 0.568 0.638 0.794 0.642 0.739
AdaMerging (Yang et al., 2023) 0.850 0.861 0.778 0.815 0.230 0.595 0.612 0.541 0.404 0.547 0.822 0.588 0.637
Consensus TA (Wang et al., 2024b) 0.892 0.894 0.866 0.882 0.376 0.596 0.714 0.665 0.520 0.630 0.881 0.664 0.715
Consensus TIES (Wang et al., 2024b) 0.898 0.893 0.847 0.862 0.314 0.631 0.689 0.604 0.461 0.630 0.862 0.644 0.695
Table 5: Multi-task accuracy of merged CLIP ViT-B/32 models on eight vision tasks.
Task SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Average Acc
Simple averaging (Wortsman et al., 2022a) 0.653 0.634 0.714 0.717 0.642 0.528 0.875 0.501 0.658
Task arithmetic (Ilharco et al., 2023) 0.552 0.549 0.667 0.789 0.802 0.697 0.973 0.504 0.692
TIES (Yadav et al., 2023) 0.598 0.586 0.707 0.797 0.862 0.721 0.983 0.542 0.725
Task arithmetic (Ilharco et al., 2023) 0.638 0.621 0.720 0.776 0.744 0.651 0.970 0.522 0.701
TIES (Yadav et al., 2023) 0.648 0.629 0.743 0.789 0.831 0.714 0.976 0.562 0.736
Fisher merging (Matena & Raffel, 2022) 0.686 0.692 0.707 0.664 0.729 0.511 0.879 0.599 0.683
RegMean (Jin et al., 2022) 0.653 0.635 0.756 0.786 0.781 0.674 0.937 0.520 0.718
AdaMerging (Yang et al., 2023) 0.645 0.681 0.792 0.938 0.870 0.919 0.975 0.591 0.801
Consensus TA (Wang et al., 2024b) 0.639 0.641 0.755 0.794 0.816 0.699 0.980 0.551 0.737
Consensus TIES (Wang et al., 2024b) 0.623 0.622 0.745 0.800 0.877 0.775 0.986 0.553 0.748
We present the per-task performance in Table 4 and Table 5. The reported normalized accuracy is computed by
Normalized $\text{Normalized accuracy}=\frac{1}{T}\sum_{t\in[T]}\frac{\text{acc}_{\text{merged}}}{\text{acc}_{\text{finetuned}}}.$
Appendix B More experiments
More on task interference. To assess our method’s effectiveness on tasks with different task similarities, we created two subsets: i) Conceptually similar subset: Composed entirely of sentiment classification tasks (SST-2, CR, MR, MPQA). ii) Conceptually dissimilar subset: Including tasks from different categories (SST-2 for sentiment classification, TREC for question classification, SUBJ for subjectivity, and MNLI for entailment). In Table 6, we report the average performance for each subset. In the similar subset, where tasks share similar skills, all merging methods perform equally well. However, in the dissimilar subset, where task skills differ and may even conflict, Localize-and-Stitch shows a significant advantage, demonstrating its ability to effectively resolve task interference.
Table 6: Performance comparison for merging similar and dissimilar tasks.
Method Average acc on similar subset Average acc on dissimilar subset
Task arithmetic 0.878 0.802
TIES 0.875 0.812
Localize-and-Stitch 0.880 0.831
@@ Appendix B More experiments
Where are the localized regions? We analyze the distribution of the localized regions for both language and vision tasks in Figure 10, both in terms of the layer index and the transformer components. For the layers, different tasks seem to occupy different layers, although the earlier layers in the network seldomly appear in the localized regions. Interestingly, most of the localized regions concentrate in the LayerNorm parameters. This pattern can possibly be attributed to a distribution shift observed in the finetuning data compared to the pretraining data, necessitating adjustments to the LayerNorm parameters to accommodate this shift.
Table 7: Single-task grafted accuracy (F1) of RoBERTa-base models on twelve NLP tasks.
Task SST-2 CR MR MPQA TREC SUBJ QNLI SNLI MNLI RTE MRPC QQP Average acc/F1
@@
Table 8: Single-task grafted accuracy of GPT2-XL models on three NLP tasks.
Task MMLU (5-shot Acc) TruthfulQA (MC2) ARC (Acc) Average
@@
Table 9: Single-task grafted accuracy of CLIP ViT-B/32 models on eight vision tasks.
Task SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Average acc
@@
Full grafted performance. We evaluate the quality of the localized regions by the grafted performance. For the $i$-th task, we only add up the localized regions in the task vectors back to the pretrained model, i.e., $\theta_{\text{pre}}+\gamma_{i}\odot\tau_{i}$. The results with a localization region of $1\%$ is shown in Tables 7, 8 and 9. For almost all tasks, using only the tiny localized region recovers nearly $99\%$ of the finetuned performance. For GPT2-XL, the performance is slightly worse because we cannot use the evaluation data for the localization step. However, the results are still strong even with surrogate datasets with similar purposes, demonstrating the flexibility and robustness of our algorithm. Overall, this shows that our localization approach is effective in locating regions containing essential skills acquired during finetuning, and the localized regions can be viewed as compact representations of the finetuned models.
@@
Runtime. We compare the runtime of Localize-and-Stitch and other baselines. We divide the algorithms into two categories: dataless and requiring data. Note that task arithmetic and TIES-Merging can fall in both categories, with the difference of whether performing hyperparameter tuning (scaling factor $\alpha$ for both, and sparsity for TIES). For the hyperparameter tuning, we follow the common practice in Ilharco et al. (2023); Yadav et al. (2023) to grid search over $\{0.1,0.2\cdots,1\}$ for the scaling factor and $\{0.1,0.2,0.3\}$ for the sparsity.
We report the runtime in Tables 11 and 11 for merging twelve NLP tasks with RoBERTa-base. For the dataless algorithms, simple averaging and task arithmetic are very efficient, as they only involve arithmetic operations on the weights. Both TIES and our dataless version requires sorting the task vectors to get the top-$k\%$ largest elements, so the runtime is slower. Compared with TIES, we do not have the step for resolving sign conflicts, so it takes less time. For algorithms requiring data, Fisher merging is the most efficient, as it uses a diagonal estimate of the Fisher information matrices with little data (256 per task). Both task arithmetic and TIES-Merging show substantial time increase, as they need to do grid search on 9 and 27 hyperparameters respectively as well as evaluating on the validation data in each run. AdaMerging takes significantly more runtime to execute compared with others, and the reason could be that entropy minimization converges slowly, as we observe that AdaMerging requires around 500 epochs to converge. Compared with other algorithms, Localize-and-Stitch executes in a relatively short amount of time, showing its effectiveness.
Appendix C Comparison between Dataless Localize-and-Stitch and TIES-Merging
Due to the similarity of the dataless version of our approach with TIES-Merging (Yadav et al., 2023), we compare them in detail.
Table 12: Accuracy (F1) comparison of Dataless Localize-and-Stitch and TIES on language tasks.
Task SST-2 CR MR MPQA TREC SUBJ QNLI SNLI MNLI RTE MRPC QQP Average acc/F1
@@ Appendix C Comparison between Dataless Localize-and-Stitch and TIES-Merging
Table 13: Accuracy comparison of Dataless Localize-and-Stitch and TIES on vision tasks.
Task SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Average acc
@@
The first step of both algorithms is similar: select the top-$k\%$ largest positions in the task vector. The primary difference lies in the selection threshold: Dataless Localize-and-Stitch selects the top-$5\%$, while TIES selects the top-$20\%$. However, as shown in Section 3.1, we find that when a localized region already contains sufficient task-specific knowledge, including more parameters only introduces more task interference. This observation could partially explain our superior performance. However, this is not the only limitation in TIES, as reducing the threshold in TIES to be $5\%$ does not yield an improved performance as demonstrated in Tables 12 and 13.
@@
Appendix D Comparison between Localize-and-Stitch and Consensus Merging
Given the similarity of our approach with Consensus Merging (Wang et al., 2024b), we compare them in detail. In Table 2, we have demonstrated Localize-and-Stitch has superior performance, and we provide further empirical analysis as follows.
Localization area.  Our method is able to localize as little as 1% of total parameters, compared to 20% required by (Wang et al., 2024b). Our small localized regions lead to less task interference, better multi-task performance and much better compression rates. In Consensus Merging, the sparsity is controlled by the hyperparameter $\lambda$, which is chosen among $\{0.2,0.3,0.4,0.5,0.6\}$. Note that is not directly used as the sparsity, rather, it serves as a scaling factor to determine the magnitude threshold for localization, as shown in Equation (5) in (Wang et al., 2024b). Intuitively, a larger imposes a stricter threshold, resulting in a smaller localized region. We present the resulting average sparsity corresponding to each choice of in Table 14. While the resulting sparsity might vary depending on the specific tasks and models, we find that all choices of $\lambda$ lead to an average sparsity of at least 25%, which is substantially larger than the 1% sparsity achieved by our localization method.
Table 14: Correspondence between $\lambda$ and sparsity.
$\lambda$ Average sparsity (language) Average sparsity (vision)
0.2 0.4977 0.5071
0.3 0.4316 0.4144
0.4 0.3864 0.3456
0.5 0.3544 0.2939
Ablation on sparsity. Similar to Appendix C, we provide an ablation study to demonstrate that the superior performance of our method is not only due to the sparsity, but also the effectiveness of our localization approach. To achieve the same sparsity level of 1% as our method, we find that we need to set the sparsity hyperparameter $\lambda$ in (Wang et al., 2024b) to be at least 0.95. This value falls significantly outside the range of the hyperparameter space they consider, indicating that TALL masks are unlikely to perform well when used to identify regions this small. Indeed, our experiments in Table 15 confirm that TALL masks with 1% sparsity yield poor performance on both vision and language benchmarks. This further illustrates the effectiveness of our localization approach in identifying tiny regions containing essential information, contributing to reduction of task interference and substantially better compression rates.
Table 15: Comparison with (Wang et al., 2024b) at the same sparsity level.
Method Sparsity Average language acc Average vision acc
Consensus TA 1% 0.477 0.518
Consensus TIES 1% 0.463 0.524
Dataless Localize-and-Stitch 5% 0.734 0.740
Localize-and-Stitch 1% 0.759 0.799
Runtime comparison. The larger runtime observed (Table 11) for Consensus TA compared to our method is primarily due to its extensive hyperparameter tuning process. Take merging vision models as an example, for the formulation of (Wang et al., 2024b), the hyperparameter tuning process is as follows: i) Tune sparsity for all tasks, which requires 40 evaluations for 5 choices of on 8 tasks. ii) Tune scaling factor , which requires 88 evaluations for 11 choices of on 8 tasks. In total, there are 128 task-specific evaluations required to select the two hyperparameters following Appendix A.2 in (Wang et al., 2024b), which we find very time-consuming in practice. In contrast, our algorithm does not require hyperparameter tuning. It operates by training on 8-shot data for 10 epochs, a process we find highly efficient.
Furthermore, hyperparameter tuning, particularly for scaling factors, has been shown to greatly influence model merging performance, making methods like Consensus TA sensitive to these choices as highlighted in (Tam et al., 2024). On the other hand, our method avoids these complexities, offering a more efficient and robust approach that enhances performance consistency.
Appendix E Details on localization
Here, we detail the skill attribution method in Panigrahi et al. (2023) and explain the difference with our formulation. Panigrahi et al. (2023) aims to localize task-specific skills contained in finetuned language models. They introduce model grafting, where for given pretrained and finetuned model parameters $\theta_{\text{pre}}$ and $\theta_{\text{ft}}$, they graft parameters of $\theta_{\text{ft}}$ in the region $\gamma$ onto the pretrained model as